How do I translate and simplify the following phrase? The difference between sixteen and one, divided by the sum of two and three.
Answer by stanbon(57219)

The difference between sixteen and one, divided by the sum of two and three.
-----------------------------
It could be (16-1)/(2+3) = 15/5 = 3.
==================================
Cheers,
Stan H.